Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates:

Machine-hours required to support estimated production 157,000
Fixed manufacturing overhead cost $ 652,000
Variable manufacturing overhead cost per machine-hour $ 4.40

Required:

1. Compute the plantwide predetermined overhead rate.

2. During the year, Job 400 was started and completed. The following information was available with respect to this job:

Direct materials $ 380
Direct labor cost $ 230
Machine-hours used 32

Compute the total manufacturing cost assigned to Job 400.

3. If Job 400 includes 60 units, what is the unit product cost for this job?

4. If Moody uses a markup percentage of 130% of its total manufacturing cost, then what selling price per unit would it have established for Job 400?
Answer #1

1.

Plantwide Predetermined overhead rate = Variable overhead cost per machine-hour + Fixed overhead cost/Estimated machine hours

= 4.40 + 652,000/157,000

= 4.40 + 4.15

= $8.55 per machine hour

2.

Overhead applied = Actual machine hours used x  Predetermined overhead rate

= 32 x 8.55

= $274

Job cost sheet

Direct material

380

Direct labor

230

Overhead applied

274

Total manufacturing cost

$884

3.

Cost per unit = Total cost/Number of units

= 884/60

= $14.73

4.

Markup percentage = 130% of total job cost

Mark up = 14.73 x 130%

= $19.15

Selling price per unit = Cost per unit + Mark up

= 14.73 + 19.15

= $33.88

Exact answer may slightly differ due to rounding off.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.